# Algorithmic idealism: what should you believe to experience next?

### v1

@@ VII Conclusions
The mere possibility (even if considered unlikely) that some aspects of Algorithmic Idealism are closer to the truth than our standard view hints at a surprising lack of imagination in the way we look at some essential questions of humanity. Humanity should colonize the galaxy in order to make a large number of conscious lives possible that would not otherwise have existed [54]? This assumes a particular metaphysics of “one single world among infinitely many possible ones” that has already been criticized by Lewis [52] and that is rejected by Algorithmic Idealism. Shutting down a computer simulation terminates its simulated conscious beings [46]? Algorithmic Idealism claims that this is generally not the case, and that the answer will depend on how much information from the outside world is fed into the simulation. But if we do not understand what to say about the simulation of agents, how can we be so sure we understand death (recall Section I)? The mere reason why we consider the view of transitioning into an infinite-duration unconscious “muted state” as the only scientifically plausible possibility may be that we have so far only seen alternatives promoting esoteric nonsense and religious wishful thinking, and that we want to signal how serious and trustworthy we are by not even addressing the question.
@@ Acknowledgments
I am grateful to Kelvin McQueen for many helpful and stimulating discussions, and in particular for his earlier contributions to what has become Section VI of the present paper.

### v2

@@ VIII Conclusions
The mere possibility (even if considered unlikely) that some aspects of algorithmic idealism are closer to the truth than our standard view hints at a surprising lack of imagination in the way we look at some essential questions of humanity. Should humanity colonize the galaxy in order to make a large number of conscious lives possible that would not otherwise have existed [65]? Arguments in favor of this typically implicitly assume a particular metaphysics of “one single real world among infinitely many possible ones” that has already been criticized by Lewis [63] and that is rejected by algorithmic idealism. Does shutting down a computer simulation terminate its simulated conscious beings [56]? This usually seems to be taken as obviously true, but algorithmic idealism claims that this is not generally the case, and that the answer will depend on how much information from the outside world is fed into the simulation. But if we do not understand what to say about the simulation of agents, how can we be so sure we understand death (recall Section I)? The mere reason why we consider the view of transitioning into an infinite-duration unconscious “muted state” as the only scientifically plausible possibility may be that we have so far only seen alternatives promoting esoteric nonsense and religious wishful thinking, and that we want to signal how serious and trustworthy we are by not even addressing the question.
These examples also demonstrate that algorithmic idealism is not merely a theory of how agents can best predict their future observations, but a radically new approach to physical reality. If it were simply some sort of theory of epistemology, then what will actually happen to an agent would be determined by the (in general unknown) physical world in which the agent would operate. In contrast, algorithmic idealism denies that the agent is fundamentally embedded into some world, and grounds whatever will happen to the agent on its self state alone. This counterintuitive starting point allows it to formulate a law (Postulate 2) which is meant to express how an agent’s fate is determined in mundane but also truly exotic situations, such as simulation and multiplication scenarios, and it makes very counterintuitive predictions (such as the possibility that agents may “fall out” of the world into which they are effectively approximately embedded, or the notion of probabilistic changelings). These predictions would not only be impossible to obtain in an epistemic approach (because, as we discussed via Restriction A, third-person claims about the world have nothing to say about first-person chances in, say, duplication scenarios), but they outright contradict the assumption of an agent that is embedded into some given but unknown world. This also indicates a clear separation from operational theories [57] as they are often formulated in the quantum information theory context: while these theories also often feature agents and their observations (such as Alice and Bob in a communication scenario), these agents are assumed to be embedded into a joint physical system (external world) that mediates their communication and observations. In algorithmic idealism, on the other hand, this external world is not postulated a priori, but an emergent statistical phenomenon, and so is the intuitive idea that Alice and Bob share the same physical world.
In summary, algorithmic idealism was not constructed with the intention to say anything directly about reality. However, the methodological starting point to ignore the usual realist assumptions led to a formulation which indeed admitted novel predictions that are not otherwise available, but at the same time made it incompatible with the received view of physical reality, and thus distinct from existing epistemic or operational approaches.
@@ Acknowledgments
I am grateful to Kelvin McQueen for many helpful and stimulating discussions, and in particular for his earlier contributions to what has become Section VII of the present paper. I would also like to thank Michael Cuffaro for helpful comments on an earlier draft.